OpenStudy (owlet):

Linear Approximation help! I need a step-by-step guidance. Let f(x)=sin(x/2).If x=1 and dx=1/2, what is dy?

OpenStudy (owlet):

what I did so far: dy/dx = f'(x) so dy = f'(x) dx dy = 1/2 cos (x/2) dx sub in values: dy = 1/2 cos (1/2) x (1/2) dy is approx 0.25 but the book says 0.22, how to get exactly 0.22?

OpenStudy (owlet):

@zepdrix @Directrix @dan815

OpenStudy (dan815):

what is this ".Ifx=1 "

OpenStudy (owlet):

there should be a space , sorry "If x=1 and dx=1/2, what is dy? "

OpenStudy (dan815):

http://prntscr.com/8zjg0f

OpenStudy (dan815):

cos(1/2)=~0.88 1/4*(cos(1/2))=~0.22

OpenStudy (amistre64):

yeah, i was gonna say use radians instead of degrees

OpenStudy (owlet):

oh I see, so my solution is right? I only have to change it to radians... thank you ! :)

OpenStudy (dan815):

=]